Miller Manufacturing has a target debt–equity ratio of .70. Its cost of equity is 14 percent, and its cost of debt is 7 percent. If the tax rate is 38 percent, what is the company’s WACC?
Answer #1

Here we need to use the debt–equity ratio to calculate the WACC. Doing so, we find:

RWACC = .14(1 / 1.70) + .07(.70 / 1.70)(1 – .38)

RWACC = .1002, or 10.02%
Answer #2

To calculate the Weighted Average Cost of Capital (WACC), we need to consider the weights of equity and debt in the company's capital structure. The WACC formula is:

WACC = (E/V) * Re + (D/V) * Rd * (1 - T)

Where: E = Market value of equity V = Total market value of the firm (E + D) Re = Cost of equity D = Market value of debt Rd = Cost of debt T = Tax rate

Given information: Target debt-equity ratio = 0.70 Cost of equity (Re) = 14% (0.14 as a decimal) Cost of debt (Rd) = 7% (0.07 as a decimal) Tax rate (T) = 38% (0.38 as a decimal)

Let's assume the total market value of the firm (V) is $100.

Now, we can calculate the market values of equity (E) and debt (D) using the target debt-equity ratio:

E = 0.70 * V = 0.70 * $100 = $70 D = (1 - 0.70) * V = 0.30 * $100 = $30

Now, we can calculate the WACC:

WACC = (E/V) * Re + (D/V) * Rd * (1 - T) WACC = ($70/$100) * 0.14 + ($30/$100) * 0.07 * (1 - 0.38) WACC = 0.70 * 0.14 + 0.30 * 0.07 * 0.62 WACC = 0.098 + 0.0133 WACC = 0.1113 or 11.13%

The company's WACC is approximately 11.13%.

answered by: Hydra Master